1102: SMTP server offline for half an hour
Both processor fans have failed and need to be replaced.
CPK messages are initially sent to the CPK mailing list, you can (un)subscribe via this link. You can also follow the service interruption messages via RSS using the link in the title under the RSS icon. If the CPK takes more time to resolve, any updates are published on this website.
For RU wide service interruption see meldingen.ru.nl.
Both processor fans have failed and need to be replaced.
The fileserver had problems, which were resolved after a server reboot. Affected network shares: acfiles ARK-Backup Bargerveen Bargerveen-Gebruikers beevee bioniccell bio-orgchem B-Ware-Arnica B-Ware-Dotter B-Ware-Erica B-Ware-Knolrus B-Ware-Lobelia B-Ware-Stratiotes celbio1 celbio2 csgi-archief csg-staf cvi desda ds ehef1 evsf2dataoud evsf2schijf1 femtospin Floron geminstr2 giphouse gissig gmi hfml-backup hfml-backup2 highres impuls introcie isis janvanhest kaartenbak kangoeroe microbiology microbiology2 milkun2 milkun3 milkun5 milkun6 milkun7 milkun7rw molbiol mwstudiereis neuroinf neuroinf2 nfstest nwibackup olympus puc RandomWalkModel Ravon-Algemeen Ravon-Foto Ravon-Projecten secres2 sigma sigmaalmanak sigmacies sigmaexchange sigmasymposium splex spmdata3 spmdata4 spmdata5 spmdata6 spmdata7 stroom thalia ucm vsc vsc1 vsc10 vsc11 vsc12 vsc13 vsc14 vsc15 vsc2 vsc3 vsc4 vsc5 vsc6 vsc7 vsc8 vsc9 vscadmin wiskalg wiskunde ...
While administering the network switchports for the move of ISIS to Mercator1 floor 0 to 3, ISC network management erroneously also changed the configuration of the network switchports of floor 4. When this was reported this morning, it could be readily remedied.
One of the mirrored disks failed. We didn’t manage to rebuild the raid set with a new disk. The system is reinstalled with two new disks. After that, data from /scratch could be recovered from one of the old disks.
The network drives will be migrated to a new server. During the migration, the network drives will be unavailable. Affected shares: acfiles, bioniccell, bio-orgchem, celbio1, CSGI-Archief, Csg-Staf, ds, evsf2schijf1, giphouse, gissig, gmi, hfml-backup, hfml-backup2, highres, impuls, introcie, isis, kaartenbak, microbiology, microbiology2, microbiologyftp, milkun2, milkun3, milkun5, milkun6, milkun7, milkun7rw, mwstudiereis, neuroinf, olympus, permissies, puc, secres2, sigma, sigmaalmanak, sigmacies, sigmaexchange, sigmasymposium, Soheto, splex, spmdata3, spmdata4, spmdata5, spmdata6, spmdata7, stroom, thalia, ucm, vsc, vsc1, vsc2, vsc3, vsc4, vsc5, vsc6, vsc7, vsc8, vsc9, vsc10, vsc11, vsc12, vsc13, vsc14, vsc15,wiskalg, wiskunde ...
The planned move of network drives from an old to a new server caused a lot of “stale NFS file handles”, which made it necessary to reboot some login servers.
The IMAP server got overloaded again. Restarting the IMAP service was required to bring it back in operation. We think it is related to snapshots and therefore regularly remove invalid snapshots from now on.
The IMAP server got overloaded. Restarting the IMAP service was required to bring it back in operation.
The server had a problem, which had started at the Monday morning reboot. A reboot of the machine solved the problem. Soon all data will be moved to a new server. This will not cause much problems for users, because everybody uses the aliases something-srv.science.ru.nl.
At Monday May 19, a major upgrade of the wireless infrastructure and its maintenance tool (called Prime) will take place by the ISC. Therefore the wireless service (i.e. Eduroam) will be interrupted several times from 8:00 pm. Existing wireless connections will be lost and new connections will not be possible for some time. Due to the complexity of this operation this maintenance will possibly be continued at Monday, May 26, again with interruptions from 8:00 pm. ...